1dm³ of 2M CH3COOH is mixed with 1dm³of 3 M etanol to form ester. The decrease in the initial rate of each solution is diluted with an equal volume of water would be:

Grade:12th pass

1 Answers

Arun
5 years ago
r = k*[CH₃COOH]*[C₂H₅OH] . 

Initial rate: r₁ = k*(a)*(b) 

The "final", r₂ =k*(a/2)*(b/2) 

(r₂/r₁) = [k(a/2)(b/2)]/ [k(a*b)] = [ab/4] / [ab] = 1/4 ⟶ r₂/r₁ = 0.25 ⟶ r₂ = (0.25)*r₁ 

The final rate, is 0.25 times the inital rate, which corresponds to answer choice: